OpenStudy (anonymous):

How old are they? Will is w years old. Ben is 3 years older. Jan is twice as old as Ben If you add their ages together the total comes to 41years old. From an equation and solve it to work out how old Will, Ben, and Jan are. Will...........years old Ben...........years old Jan............years old In how many years time will Jan be as old as Will? in................years time Explain how you figured it out. ............................................................................. SHOW YOUR WORK!

OpenStudy (texaschic101):

Will = w Ben = w + 3 Jan = 2(w + 3) w + (w + 3) + 2(w + 3) = 41 can you solve it now ?

OpenStudy (anonymous):

yesss thank you soooooo much!